Data from n = 272 eruptions of the Old Faithful geyser in the United States are used to study the straight-line relationship between x = time to next eruption (minutes) and y = duration of the present eruption (minutes).

95% CI for β0: (-2.189, -1.559), 95% CI for β1: (0.071, 0.080).

(a) Using the above information, compute estimates for β0 and β1
Answer #1

a) as we know that estimated mean is at center of confidence interval

therefore

estimate for β0 =(-1.559+(-2.189))/2=-1.874

estimate for β1 =(0.071+0.080)/2=0.0755